OpenStudy (anonymous):

help simplify -7(x-2)+5(3-x)-4x. Show steps too please! xO

OpenStudy (anonymous):

okay you start by multiplying the numbers outside of the brackets into the brackets. Try to do this yourself, post your result and I will let you know if you did something wrong.

OpenStudy (anonymous):

okay! give me a moment to do it! xD thank u for helping!

OpenStudy (anonymous):

-7(x-2)+5(3-x)-4x -7(-2x)+5(-3x)-4x -I think that's not right. X^X ugh is it right from here? am I doing it right or wrong.?

OpenStudy (anonymous):

. . . I dont think you understood me so I will show you how to do it for -7(x-2) BTW I am reading the inside term as x minus two -7(x-2)=-7*x-(-7)*2=-7x+14 where * is the multiplication sign

OpenStudy (anonymous):

uhm.. u lost me. )''X

OpenStudy (anonymous):

oh wait I get what u mean!

OpenStudy (anonymous):

but don't I have to do that all together? That's what confuses me, I get that part but what do I do with the +5? Do I combine that with the (3-x)? Or -4x?

OpenStudy (anonymous):

the +5 is confusing me with it all.

OpenStudy (anonymous):

The multiplication only applies to the terms inside the brackets so you would only multiply the term (3-x) by five since that is the term inside the brackets.

OpenStudy (anonymous):

u mean like what u did with the -7. okay... I think I got. can I try it all together and can u check it? It'll take a moment. X''(

OpenStudy (anonymous):

Of course! I already have the solution I got ready.

OpenStudy (anonymous):

wah...I haven't even...X^X ..figured it out. <:'') You make it sound easy.

OpenStudy (anonymous):

Can you tell me how you did it and explain it all? TT^TT im stuck again at the beginning. im not cheating! im just trying to understand it by observation. please...

OpenStudy (anonymous):

I'm in university, and a math major so this stuff is a little on the easy side for me. I can start to tell you about some math concepts which are quite interesting.

OpenStudy (anonymous):

Did you get the first step I told explained to you?

OpenStudy (anonymous):

X^X no. I don't ...half get it.

OpenStudy (anonymous):

x''O to tell u the truth I have state test tomorrow and I slacked off all this week from studying. Now, i'm last minute studying, hoping I remember it all. but I have terrible memory and I forgot everything...

OpenStudy (anonymous):

type out the equation without multiplying anything but factoring in the numbers example: 3(x-5)=3*x-3*5 is what I want you to do for your question.

OpenStudy (anonymous):

okay! @0@

OpenStudy (anonymous):

3(x-5)=3*x-3*5= 3x-15? 0-0?

OpenStudy (anonymous):

3x-15(-4x)= 3x*(-4x)-15*(-4x) ? 0-0?

OpenStudy (anonymous):

(3x*-4)- (15*x) is this right instead?

OpenStudy (anonymous):

Why are you multiplying it by -4x? This does not change during this step you can just keep it as your last term throughout. If you want you can even separate your question into three parts for now. instead of simplifying -7(x-2)+5(3-x)-4x simplify -7(x-2) then simplify 5(3-x) then simplify -4x I'll give you a hint: the last one cannot be simplified further.

OpenStudy (anonymous):

0o0! ooohhhhh...okay.

OpenStudy (anonymous):

tell me the answer to each one of the parts I just posted and then we will combine them.

OpenStudy (anonymous):

okay well u got the first one which was: -7x+14 the second (may be wrong) :3x-15 last one: -4x right?

OpenStudy (anonymous):

remember 5(3-x) means 5*3-5*x

OpenStudy (anonymous):

yes. 0o0

OpenStudy (anonymous):

so try solving 5*3-5*x and tell me what that is

OpenStudy (anonymous):

okay 5*3=15 and 5*x= 5x?

OpenStudy (anonymous):

yes so 5*3-5*x is?

OpenStudy (anonymous):

15-5x? 0-0

OpenStudy (anonymous):

yes so 5(3-x)=15-5x

OpenStudy (anonymous):

0o0 okay!

OpenStudy (anonymous):

(-7x+14)(15-5x)(-4x) okay. 0o0

OpenStudy (anonymous):

so now we have part 1, part 2 and part 3 of your question simplified as much possible while they are separated so now lets recombine them into the original question. Why are you dropping the addition and subtraction? they do not disappear with this method

OpenStudy (anonymous):

oh. 0-0 I ..just thought to show the part 1,2 ,3 separated. TT^TT

OpenStudy (anonymous):

Put the simplified expressions back into your original question and tell me what you get, do not simplify yet.

OpenStudy (anonymous):

I'm assuming that you are no longer here and plan to leave the site shortly if you do not respond.

OpenStudy (anonymous):

I am logging off now, goodnight.

OpenStudy (anonymous):

oh thank you! something happen with my internet and it messed up! thank u for your help! xD

OpenStudy (anonymous):

goodnight! xD